Consider the information in the file named HW2 - Production Data for a Single Firm. Currently the nominal wage rate is W = $1,000 per worker and the price of the output is P = $10 per unit. We already know that the number of workers the firm will want to hire depends on the relationship between the nominal wage it pays each worker and the price per unit of the output that worker produces. This relationship is captured by the ratio of the two called the real wage.  

If the wage rate increases by 68% and the price level increases by 20%, the real wage will change to ————units, the firm will hire ————workers, and will produce a total of ————units of output.

On the other hand, if the wage rate increased by 12% and the price level increased by 40%, the real wage would change to ————units, the firm would hire ————workers, and would produce a total of ————units of output.

Total Labor Product (Persons) (Tons) 0 0 200 1 2 380 3 540 4 680 800 6 900 980 8 1040 1080 10 1100 Ln
Answer #1
Labour Total product Marginal product of labour
0 0 0
1 200 200
2 380 180
3 540 160
4 680 140
5 800 120
6 900 100
7 980 80
8 1040 60
9 1080 40
10 1100 20

Real wage = Nominal wage/ price

Real wage=$1000/$10=100

If the wage rate increases by 68% and the price level increases by 20%, the real wage will change to 140 units, the firm will hire 4 workers, and will produce a total of 680 units of output.

New Nominal wage rate= 1000(1.68)= 1680

New price of the output= 10(1.20)= 12

New real wage= 1680/12= 140

Firm will hire till real wage= Marginal product of labour

TP at 10 labours= 680

On the other hand, if the wage rate increased by 12% and the price level increased by 40%, the real wage would change to 80 units, the firm would hire 7 workers, and would produce a total of 980 units of output.

New Nominal wage rate= 1000(1.12)= 1120

New price of the output= 10(1.40)= 14

New real wage= 1120/14= 80

Firm will hire till real wage= Marginal product of labour

TP at 7 labours= 980

If it helps kindly upvote.